OpenStudy (anonymous):

aIf an object travels in a circle 32pie/5 radians, at what degree measure is it, relative to where it started?

OpenStudy (dumbcow):

just convert radians to degrees....pi = 180 degrees --> 32*180/5 = 1152 every time around is 360 360*3 = 1080 1152 - 1080 = 72 therefore its 72 degrees from where it started